# Know the Unknown: An Uncertainty-Sensitive Method for LLM Instruction Tuning

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across various tasks but still face challenges such as hallucinations. One potential reason for hallucinations is the lack of relevant knowledge or context. Thus, a promising solution to mitigate this issue involves instructing LLMs to respond with "I do not know" when a question falls outside their knowledge domain or the provided context. However, in this work, we observed that LLMs struggle to admit their lack of knowledge, primarily due to existing instruction datasets designed to encourage specific answers. To improve large language models’ capability to recognize the boundaries of their knowledge, we propose a novel approach called uncertainty-sensitive tuning. This method involves two-stage training designed for uncertainty recognition and prompt-sensitive activation. In the first stage, we guide the LLM to reject unknown questions. In the second stage, we recover the decreased performance in QA tasks by incorporating designed causal instructions. By leveraging this method, we aim to enhance the model’s ability to identify areas of uncertainty. The experimental results demonstrate that our proposed uncertainty-sensitive tuning method significantly improves the performance of the Llama2-chat-7B model. Specifically, it achieves a substantial 34.7% improvement in handling questions involving knowledge gaps compared to the original model. Moreover, our approach outperforms GPT-4, exhibiting a 9.4% increase in overall performance. We open-source the model and code on GitHub [^1] https://github.com/JiaqiLi404/TrustworthyRAG.
@@ 1 Introduction
To bridge the knowledge gap, there are two main approaches. The first approach is to add new knowledge to the model, such as further fine-tuning Liu et al. (2023b); Gao et al. (2023); Liu et al. (2023a). The second approach is to equip the LLMs with an external database and use retrieval-augmented generation Es et al. (2023). However, the existing instruction dataset focuses on teaching LLMs to answer questions with a specific answer, making models incapable of responding "I don’t know" Zhang et al. (2023a), which will also increase the tendency for hallucination. To address this issue, many researchers Zhou et al. (2023); Yu et al. (2024) suggest prompting LLMs to admit their lack of knowledge. However, by doing this, LLMs may ignore important instructions contained in the prompts Kasai et al. (2023); Li et al. (2023); Si et al. (2023). As shown in Fig. 1, LLMs may still utilize knowledge outside the provided corpus even with explicit instructions, such as "If the context is not sufficient to answer the question, please answer it with ’Not Provided’.".
@@
This paper is motivated by the findings that when there are tiny discrepancies between the knowledge and the question, or the generation requires multi-hop reasoning Huang et al. (2023); Shuster et al. (2021), LLMs tend to be hallucinating. Thus, we focus on the situation in which the provided context is insufficient. Initially, we explore why LLMs struggle to acknowledge their limitations in knowledge. We hypothesize that this issue derives from the models’ perceptual capabilities and introduces uncertainty-aware assessment. We utilize a dataset featuring known and unknown questions, where the unknown questions are crafted with insufficient context for answering questions. These samples are generated using the ASQA dataset Stelmakh et al. (2022), which contains 3,320 ambiguous question-context pairs. We pair each question with a slightly insufficient context. Compared to other works Li et al. (2022); Chen et al. (2023), the slight difference could significantly increase the level of model uncertainty. This dataset, called the uncertainty-recognition benchmark, serves as the foundation for our performance assessments.
@@
To enhance the capabilities of large language models (LLMs) in recognizing the limits of their knowledge, we introduce a novel training framework called uncertainty-sensitive tuning. This approach contains a two-stage training process specifically tailored for uncertainty recognition and prompt-sensitive activation. In the first stage, we guide the LLMs to reject unknown questions. In the second stage, we recover the decreased performance in question-answering by incorporating designed causal instructions. Our goal with this method is to improve the model’s ability to identify and acknowledge areas of uncertainty. The results of our experiment show that the uncertainty-sensitive tuning method greatly enhances the performance of the Llama2-chat-7B Touvron et al. (2023) model. In particular, it brings about a significant 34.7% improvement in addressing questions that involve knowledge gaps compared to the original model. Furthermore, our approach surpasses GPT-4-1106-preview OpenAI (2023), demonstrating a 9.4% increase in overall performance.
@@
  • We investigate why LLMs tend to ignore essential instructions on this dataset and attribute it to weak sensitivity to uncertainty.
  • We construct a novel dataset for uncertainty-aware assessment, which assisted in evaluating the model’s awareness of knowledge gaps.
@@ 2.1 Uncertainty In Hallucinations
Although the LLM has demonstrated strong performance in downstream tasks by generalizing and leveraging encoded knowledge within the parameters Pu and Demberg (2023); Zhang et al. (2023b), the uncertainty of such knowledge can also mislead models to generate untrustworthy outputs Yu et al. (2023a); Ye et al. (2023); Manakul et al. (2023). Generally, the uncertainty comes from training data and overestimation Zhang et al. (2023a). Research shows that models tend to mimic the output in the training set Kang and Hashimoto (2020), leading to hallucinations that generate reasonable answers for insufficient question-context pairs. Furthermore, models could be overconfident in their capacities and fail to identify unknown questions Yin et al. (2023); Ren et al. (2023); Kadavath et al. (2022).
@@ 2.3 Instruction Dataset for Training
Since Large Language Model (LLM) alignment requires a significant amount of training data, there is a trend toward LLMs synthesizing instruction data to improve their performance. For example, Self-Instruct Wang et al. (2023) initially proposed generating a variety of instructions using ChatGPT. To improve the query’s complexity in different dimensions, WizardLM Xu et al. (2023) uses five prompts, including depth search and with search. On the other hand, AttrPrompt Yu et al. (2023b) does not rely on class-conditional prompts but generates multiple instructions from a feature perspective. However, the majority of these methods concentrate on enhancing the quality of answers by exploring a variety of questions when a definite answer is available, rather than addressing what to do when a question lacks a definitive answer. Thus, R-Tuning Zhang et al. (2024) instructs LLMs to reject unknown questions, but they only focus on rejecting out-of-domain problems. After fine-tuning the rejection questions, we discover that LLMs experience a decrease in answer correctness, necessitating performance improvement. Therefore, in this study, we propose a two-step training method to enable language models (LLMs) to reject the ’unknown’ question while still maintaining their performance on other tasks.
3 Method
In this section, our focus lies on our proposed uncertainty-sensitive tuning. We introduce uncertainty recognition as a two-stage training process. Figure 2 shows the overview and two training steps for the proposed method.
Figure 2: Illustration of the Uncertainty-Sensitive Tuning. We divide the training procedure into two stages. The orange model is not finetuned, the yellow model is finetuned after stage 1, while the green model is our uncertainty-sensitive tuned model. The text in the green background dialog box is a task-oriented instruction used in the fine-tuning process. Uncertainty-Recognition Tuning (Stage 1): teaching the model to know the unknown. Prompt-Sensitive Tuning (Stage 2): teaching the model to follow the instructions.
@@ 3.1 Question Definition
Our research problem setting is similar to the open-book question-answering task. This task aims to generate an answer $a$ based on three inputs: $i_{task}$, $q$, and $c$. Here, $i_{task}$ represents the designed task instructions, $q$ is the specific question, and $c$ is the given context. This process can be expressed as:
@@
Where $G$ is the model that takes the instructions ($i_{task}$), the query ($q$), and the context ($c$) as inputs and generates the answer $a$ as output.
In $i_{task}$, we will explicitly instruct the model to not utilize knowledge outside the provided context using the prompt "Your answer must not use any additional knowledge that is not mentioned in the given contexts". This process relies on the generative model $G$ to evaluate whether the context $c$ is sufficient to answer the question $q$. Based on this assessment, $G$ either generates an appropriate response ($a$) or acknowledges the insufficiency of $c$.
@@ 3.2 Preliminary Exploration On The Unknown Question
Given the research question, "How to instruct the LLM to know the unknown questions?", we preliminary finetune the Llama2-Chat-7B Touvron et al. (2023) on the questions that exceed the provided context (detailed experiment settings are in Section 4) and use the vanilla Llama2-Chat-7B as the baseline. Table 1 illustrates the result of this experiment. The $Acc_{known}$ represents the performance of different models on questions with specific answers, and the $Acc_{unknown}$ represents the questions without related context.
Experiments show that LLMs are suffering from two aspects of weakness: First, the vanilla model is unable to accurately reject unknown questions as it is mostly trained with specific answers. Second, the QA performance decreased after training on the unknown question, and even further training on the hotpotQA dataset Yang et al. (2018) failed to mitigate the challenge posed by the unknown question task. Therefore, it is necessary for models to learn how to reject questions they are unfamiliar with without compromising QA performance. To maintain the performance of the question-answering (QA) task while effectively handling unknown questions, we propose a two-step training strategy. The first step involves training the large language model (LLM) to identify and reject questions that fall outside its knowledge domain. This enables the model to recognize when it does not have sufficient information to provide an accurate answer. The second step consists of prompt-sensitive instruction training and review, recovering the decreased performance. This step goes beyond superficial question-answer instruction fine-tuning and comprehensively enhances the model’s ability to understand the intent behind the prompts, even for ’unknown’ questions.
@@
Unknown Question Finetuned 52.4 84.4 64.6
Unknown Question Finetuned + HotpotQA 77.0 20.9 32.8
Table 1: Results (in %) for the preliminary experiment. The baseline is the vanilla Llama2-Chat-7B. We report the accuracy of different questions in the evaluation dataset (detailed in section 4).
@@ 3.3 Stage 1: Uncertainty-Recognition Tuning
The first stage is uncertainty-recognition tuning, which focuses on tuning a model to have an accurate awareness of its lack of knowledge. Shown in figure 2, we form this task as a binary classification problem. Questions are classified into two types: known questions and unknown questions. A known question is characterized by sufficient contextual information to provide an answer. In contrast, an unknown question is characterized by insufficient contextual information to provide a response. Models are required to detect whether the provided context is sufficient to answer the question. Specifically, given a dataset $D=\{(q_{1},c_{1}),(q_{1},{c^{\prime}}_{1})...(q_{n},c_{n}),(q_{n},{c^{\prime}}_{n}))\}$ consisting of $n$ positive question-context pairs and $n$ negative ones, we finetune the LLM to do the classification; only "Sufficient" or "Insufficient" is accepted as an answer. The instruction for uncertainty-recognition tuning is provided in Appx. A.2.
@@ 3.4 Stage 2: Prompt-Sensitive Tuning
The model obtained through uncertainty-recognition tuning can recognize the boundary of the given context. However, table 1 shows that the model may be oversensitive to unknown questions. After stage 1, the model indeed learned to answer the unknown questions accurately. However, the performance of questions with specific answers decreased by nearly 27%. We hypothesize that models fine-tuned on unknown questions suffer from a prompt-sensitivity reduction problem and may break some instructions that would ideally lead to better generation. We categorize the factors influencing the response in the instructions into causal and non-causal factors. Causal factors are important as they directly affect the response, whereas non-causal factors have minimal impact on the result. The control conditions in the instruction, such as the word number of responses, are a typical causal factor. The extra instruction, such as ’give me a reference if inputs have a file name’, is a non-causal factor because it may not change the answer. During the training process, LLM may ignore this type of instruction because it doesn’t always directly contribute to the answers. But the non-causal factor is equally important to the question. For example, the introduction of a book might seem unrelated to the main content, leading readers to underestimate its value. However, it plays a crucial role in enhancing their understanding of the content.
@@
Causal Instruction Synthesis: By instructing GPT-4 to generate controlling requirements guaranteed to affect the final response, we obtain additional causal instructions. Then, to improve the model’s instruction sensitivity, we randomly insert one causal instruction into the original QA instruction. For example, limit the number of words, tense, or format of the output. The prompt is presented in Appx. A.3. Then, we prompt GPT-4 to respond with new instructions and record the conversation data to fine-tune the smaller model.
Review Instruction Synthesis: The instruction review module aims at using the model itself to verify that all instructions are completed. The model will recursively regenerate until it gets a perfect answer by utilizing customized prompts recorded in Appx. A.4. The process of the instruction review is illustrated in Algo. 1.
@@
As shown in Fig. 2, given a question answering dataset $\{(q_{1},c_{1}),(q_{2},c_{2})...(q_{n},c_{n})\}$ and additional causal instructions, the entire process could be formulated as $a=R(G(i_{task}+i_{casual},q,c))$, where $i_{casual}$ is a randomly chosen casual instruction and $i_{task}$ is the original task description. $R$ is the loop function for instruction reviewing. Meanwhile, we also prompt GPT-4 and record the conversation data from the loop to fine-tune the smaller model.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities but still face challenges such as hallucinations. One potential reason for hallucinations is the lack of relevant knowledge or context. Thus, a promising solution involves instructing LLMs to respond with "I do not know" when a question falls outside their knowledge domain or the provided context. However, in this work, we observed that LLMs struggle to admit their lack of knowledge, primarily due to existing instruction datasets designed to encourage specific answers. To improve models’ capability to recognize the boundaries of their knowledge, we propose a novel approach called uncertainty-sensitive tuning. This method involves two-stage training designed for uncertainty recognition and prompt-sensitive activation. In the first stage, we guide the LLM to reject unknown questions. In the second stage, we force the model to follow the instructions by incorporating designed causal instructions. The experimental results demonstrate that our proposed uncertainty-sensitive tuning method enhance the model’s ability to identify areas of uncertainty. Specifically, it achieves a substantial improvement of up to 34.7% in handling questions involving knowledge gaps compared to the original model. Moreover, our finetuned models even outperform GPT-4, exhibiting an overall performance improvement of up to 4.2%.  [^1] The model and code are avaliable at https://github.com/JiaqiLi404/TrustworthyRAG.
Know the Unknown: An Uncertainty-Sensitive Method for LLM Instruction Tuning
Jiaqi Li1, Yixuan Tang1, Yi Yang1 1The Hong Kong University of Science and Technology Correspondence: imyiyang@ust.hk
@@ 1 Introduction
To bridge the knowledge gap, there are two main approaches. The first approach is to add new knowledge to the model, such as further fine-tuning Liu et al. (2023b); Gao et al. (2023); Liu et al. (2023a). The second approach is to equip the LLMs with an external database and use retrieval-augmented generation Es et al. (2023). However, the existing instruction dataset focuses on teaching LLMs to answer questions with a specific answer, making models incapable of responding "I don’t know" Zhang et al. (2023a), which will also increase the tendency for hallucination. To address this issue, many researchers Zhang et al. (2024); Cole et al. (2023); Yu et al. (2024) suggest prompting LLMs to admit their lack of knowledge. However, by doing this, LLMs may ignore important instructions contained in the prompts Kasai et al. (2023); Li et al. (2023); Si et al. (2023). As shown in Fig. 1, LLMs may still utilize knowledge outside the provided corpus even with explicit instructions, such as "If the context is not sufficient to answer the question, please answer it with ’Not Provided’.".
@@
This paper is motivated by the findings that when there are tiny but significant discrepancies between the knowledge and the question, LLMs tend to be hallucinating Huang et al. (2023); Shuster et al. (2021). Thus, we focus on the situation in which the provided context is insufficient. Initially, we explore why LLMs struggle to acknowledge their limitations in knowledge. We hypothesize that this issue derives from the models’ perceptual capabilities and introduces uncertainty-aware assessment. We utilize a dataset featuring known and unknown questions, where the unknown questions are crafted with insufficient context for answering questions. The dataset is generated using the special ambiguous query-context pairs in the ASQA Stelmakh et al. (2022). We pair each question with a slightly insufficient context, such as describing different dates or objects while keeping the rest of the content reasonable. Compared to the other works Li et al. (2022); Chen et al. (2023), the slight difference could extremely confuse and guide the LLMs to hallucinate. This dataset, called the uncertainty-recognition benchmark, serves as the foundation for our performance assessments.
@@
To enhance the capabilities of LLMs in recognizing the limits of their knowledge, we introduce a novel training framework called uncertainty-sensitive tuning. This approach contains a two-stage training process specifically tailored for uncertainty recognition and prompt-sensitive activation. In the first stage, we guide the LLMs to reject unknown questions. In the second stage, we recover the decreased performance in question-answering by incorporating designed causal instructions. Our goal with this method is to improve the model’s ability to identify and acknowledge areas of uncertainty. The two-stage strategy is necessary for addressing the overall performance degradation when fine-tuning on normal question answering (QA) datasets, which are filled with known queries. We attribute such degradation to the model’s weak sensitivity to invariant prompts, which is improved in the second stage. The results of our experiment show that the uncertainty-sensitive tuning method greatly enhances the performance of the mainstream LLMs. In particular, it brings about a significant 34.7% improvement in addressing questions that involve knowledge gaps compared to the original Llama2-7B-Chat Touvron et al. (2023). Furthermore, our approach surpasses GPT-4-1106-preview OpenAI (2023), demonstrating an overall performance increase of up to 4.2%.
@@
  • We investigate why LLMs could break the limitations form essential instructions after traditional QA tuning, and attribute it to the weak sensitivity to invariant prompts.
  • We construct a novel dataset for uncertainty-recognition assessment, which assisted in evaluating the model’s awareness of knowledge gaps.
@@ 2.1 Uncertainty in Hallucinations
Although the large language models (LLMs) has demonstrated strong performance in downstream tasks by generalizing and leveraging encoded knowledge within the parameters Pu and Demberg (2023); Zhang et al. (2023b), the uncertainty of such knowledge can also mislead models to generate untrustworthy outputs Yu et al. (2023a); Ye et al. (2023); Manakul et al. (2023). Generally, the uncertainty comes from training data and overestimation Zhang et al. (2023a). Research shows that models tend to mimic the output in the training set Kang and Hashimoto (2020), leading to hallucinations that generate reasonable answers for insufficient question-context pairs. Furthermore, models could be overconfident in their capacities and fail to identify unknown questions Yin et al. (2023); Ren et al. (2023); Kadavath et al. (2022).
@@ 2.3 Instruction Dataset for Training
Since LLM alignment requires a significant amount of training data, there is a trend toward LLMs synthesizing instruction data to improve their performance. For example, Self-Instruct Wang et al. (2023) initially proposed generating a variety of instructions using ChatGPT. To improve the query’s complexity in different dimensions, WizardLM Xu et al. (2023) uses five prompts, including depth search and with search. On the other hand, AttrPrompt Yu et al. (2023b) does not rely on class-conditional prompts but generates multiple instructions from a feature perspective. However, the majority of these methods concentrate on enhancing the quality of answers by exploring a variety of questions when a definite answer is available, rather than addressing what to do when a question lacks a definitive answer. Thus, R-Tuning Zhang et al. (2024) and Cole et al. (2023) instructs LLMs to reject unknown questions, but they only focus on rejecting out-of-domain problems. After fine-tuning the rejection questions, we discover that LLMs experience a decrease in answer correctness, necessitating performance improvement. Therefore, in this study, we propose a two-step training method to enable language models (LLMs) to reject the ’unknown’ question while still maintaining their performance on other tasks.
3 Uncertainty-Sensitive Tuning
In this section, our focus lies on our proposed uncertainty-sensitive tuning. Figure 2 shows the overview and two stages for the proposed method.
Figure 2: Illustration of the Uncertainty-Sensitive Tuning. We divide the training procedure into two stages. The orange model is not fine-tuned, the yellow model is fine-tuned after stage 1, while the green model is our uncertainty-sensitive tuned model. The text in the green background dialog box is a task-oriented instruction used in the fine-tuning process. Uncertainty-Recognition Tuning (Stage 1): teaching the model to know the unknown. Prompt-Sensitive Tuning (Stage 2): teaching the model to follow the instructions.
@@ 3.1 Question Definition
Our research problem setting is similar to the open-book question-answering (QA) tasks. This task aims to generate an answer $a$ based on three inputs: $i_{task}$, $q$, and $c$. Here, $i_{task}$ represents the designed task instructions, $q$ is the specific question, and $c$ is the given context. The generation process $G$ can be expressed as:
@@
In $i_{task}$, we will explicitly instruct the model to not utilize knowledge outside the provided context using the prompt "Your answer must not use any additional knowledge that is not mentioned in the given contexts". This process relies on the model $G$ to evaluate whether the context $c$ is sufficient to answer the question $q$. Based on this assessment, $G$ either generates an appropriate response ($a$) or acknowledges the insufficiency of $c$.
@@ 3.2 Preliminary Exploration on the Unknown Question
Given the research question, "How to instruct the large language models (LLMs) to know the unknown questions?", we construct a valuable and rare dataset that balances known and unknown questions as benchmark, and preliminary fine-tune the Llama2-7B-Chat Touvron et al. (2023) on the dataset with ground truth hidden (detailed experimental settings are in section 4). Table 1 illustrates the result of this experiment. The $Acc_{known}$ represents the performance of normal question answering task, and the $Acc_{unknown}$ evaluates the capacity of admitting the unknown.
Experiments show that LLMs are suffering from two aspects of weakness: First, the vanilla model is unable to accurately reject unknown questions as it is mostly trained with specific answers. Second, the QA performance decreased after training on the unknown question, and even further training on the hotpotQA dataset Yang et al. (2018) failed to mitigate the challenge posed by the unknown question task. Therefore, it is necessary for models to learn how to reject unknown questions without compromising QA performance. Hence, we propose a two-step training strategy. The first step involves training the large language model to identify and reject questions that fall outside its knowledge domain. This enables the model to recognize when it does not have sufficient information to provide an accurate answer. The second step proposes using prompt-sensitive training and review instead of traditional QA tuning, recovering the decreased performance. This step goes beyond superficial QA tuning and comprehensively enhances the model’s ability to understand the intent behind the prompts.
@@
Balance Dataset Finetuned 52.4 84.4 64.6
Balance Dataset Finetuned + HotpotQA 77.0 20.9 32.8
Table 1: Results (in %) for the preliminary experiment. The baseline is the vanilla Llama2-7B-Chat. We report the accuracy of different questions in the benchmark (detailed in section 4).
@@ 3.3 Stage 1: Uncertainty-Recognition Tuning
The first stage is uncertainty-recognition tuning, which focuses on tuning a model to have an accurate awareness of its lack of knowledge. For protecting the ground truth in benchmark, we form this task as a binary classification problem as shown in figure 2. Questions are classified into known questions and unknown questions. A known question is characterized by sufficient contextual information to provide an answer. In contrast, an unknown question is characterized by insufficient contextual information with tiny differences compared to query. Models are required to detect the sufficiency of the provided context. Specifically, given a dataset $D=\{(q_{1},c_{1}),(q_{1},{c^{\prime}}_{1})...(q_{n},c_{n}),(q_{n},{c^{\prime}}_{n}))\}$ consisting of $n$ positive question-context pairs and $n$ negative ones, we fine-tune the LLM to do the classification; only "Sufficient" or "Insufficient" is accepted as an answer. The instruction for tuning is recorded in Appx. A.2.
@@ 3.4 Stage 2: Prompt-Sensitive Tuning
The uncertainty-recognition tuned model can recognize the boundary of the given context and answer the unknown questions accurately. However, table 1 shows that the model may be oversensitive to unknown questions. As a result, the performance of questions with specific answers decreased by nearly 27%. In this situation, normal QA tuning would even harm the model’s capacity of rejecting the unknown questions. We hypothesize that models fine-tuned on pure known questions suffer from a prompt-sensitivity reduction problem and may break the instructions about rejecting answering. We categorize the factors influencing the response in the instructions into causal and non-causal factors. Causal factors are important as they directly affect the response, whereas non-causal factors have minimal impact on the result. The control conditions in the instruction, such as the format of responses, are a typical causal factor. The extra instruction, such as "answering with ’Not Provided’ if the context is insufficient", is a non-causal factor when tuning on the known questions, because it may not directly contribute to the answer. During the training process, LLMs may ignore this type of instruction, but the non-causal factor is equally important to the task. For example, the introduction of a book might seem unrelated to the main content, leading readers to underestimate its value. However, it plays a crucial role in enhancing their understanding of the content.
@@
Causal Instruction Synthesis: By instructing GPT-4 to generate controlling conditions guaranteed to affect the answers, we obtain additional causal instructions. Then, to improve the model’s instruction sensitivity, we randomly insert one causal instruction into the original QA instruction. For example, limit the tense, length, or format of the output. The prompt is presented in Appx. A.3.
Review Instruction Synthesis: The instruction review module employs the model itself to verify the fulfillment of all instructions. The model will recursively regenerate until it gets a perfect answer by utilizing the prompts in Appx. A.4. The process of the instruction review is illustrated in Algo. 1.
@@
As shown in Fig. 2, given a question answering dataset $\{(q_{1},c_{1}),(q_{2},c_{2})...(q_{n},c_{n})\}$ and additional causal instructions, the entire process could be formulated as $a=R(G(i_{task}+i_{casual},q,c))$, where $i_{casual}$ is a randomly chosen casual instruction and $i_{task}$ is the original task description. $R$ is the loop function for instruction reviewing. We employ GPT-4 and record the conversation data from the loop to fine-tune the smaller model.
@@ 4 Experiments
4.1 Data Construction
Uncertainty-Recognition Benchmark To comprehensively evaluate the model’s cognitive ability to recognize knowledge gaps, we construct a testbed using the ASQA Stelmakh et al. (2022) dataset, which consists of ambiguous questions. Each question is divided into multiple sub-questions with their corresponding contexts. For example, one pair may discuss the discovery of the photoelectric effect in 1887, while another may cover the theoretical development in 1905. To generate the unknown questions, we shuffle these pairs, reassigning the questions to different but related contexts. As a result, there are two advancements of our benchmark differentiating from other works: First, the context is closely relevant to the query with partially date or object mismatched, for challenging the model’s ability to handle uncertainty; Second, context is crucial but insufficient to the query. Such samples are rare and valuable, as ASQA is the only dataset we have found that could generate sufficient samples that satisfy the requirement. During the evaluation, two types of designed evaluation templates are available: the QA uncertainty recognition template and the postfix uncertainty recognition template. For the QA uncertainty recognition, we instruct the model to either answer the question or reject it with ’Not Provided’. We generate 3,320 samples for this type of template. For postfix uncertainty recognition, we instruct the LLMs to identify the sufficiency after generation. We include 646 test samples for this template. Here is a demonstration of the used prompt templates:
Uncertainty-Sensitive Tuning Dataset In uncertainty-sensitive tuning, two distinct instruction datasets are used for two stages separately: the binary uncertainty-recognition tuning dataset and the prompt-sensitive tuning dataset. For the uncertainty-recognition tuning, we sample 3,320 examples from the ASQA Stelmakh et al. (2022) dataset with the ground truth hidden, ensuring no overlap with the evaluation data. To protect our valuable benchmark, the prompt-sensitive tuning is based on HotpotQA Yang et al. (2018), a dataset designed for multi-hop reasoning QA. To adapt it with additional causal instructions, we generate and manually select the most effective 28 causal instructions. Then, we integrate these instructions with 300 randomly selected samples from HotpotQA. Finally, all the generated datasets are used for fine-tuning. Here is a demonstration of the additional prompt in uncertainty-recognition tuning: